# IFT-UAM/CSIC-23-127

### v1

@@ Introduction
The fact that an action for the higher-rank forms must also contain the lower-rank ones can be turned into an advantage if one manages to give consistence to the simultaneous presence of dual fields in the action. A solution to this problem is to use extensions of the Pasti-Sorokin-Tonin formalism [18] which is based on the introduction of an auxiliary scalar field in the action. This method has been used in Ref. [19, 20] to construct covariant actions of $\mathcal{N}=2B,d=10$ supergravity containing the 4-form but also 8-form duals of the scalar fields and also in Ref. [21] to construct an action of $\mathcal{D}=1,d=11$ supergravity containing the 3-form and its dual 6-form simultaneously. By dimensional reduction one can obtain an action of $\mathcal{N}=2B,d=10$ supergravity containing the fundamental and dual fields [22]. In a slightly different context, it has been used in Ref. [23] to construct a covariant worldvolume action of the M5-brane.
An alternative solution, proposed in Ref. [24], consists in including all the fields and treating them all on an equal footing as independent (making it democratic). This procedure doubles the degrees of freedom and one has to impose by hand the twisted duality [25, 26] relations between $(d+1)$- and $(d-p-3)$-forms only after the equations of motion have been derived from the action. Since the duality relations are not derived from the action, one is actually dealing with a pseudoaction. The pseudoaction introduced in Ref. [27] for $\mathcal{N}=2B,d=10$ supergravity, which includes the 4-form with selfdual 5-form field strength provides a good example: it contains an unconstrained (not selfdual) 4-form which describes twice the degrees of freedom of the selfdual one and the selfduality constraint must be imposed after the equations of motion have been derived.
@@
When the theory contains several scalar fields which parametrize a non-linear $\sigma$-model, things become more complicated. The shift symmetries are isometries of the $\sigma$-model metric. One can always use coordinates (scalar fields) adapted to a given isometry. In those coordinates the scalar field shifted by the isometry does not occur explicitly in the metric and the action can always be written in terms of its derivatives only. The equation of motion will be a total derivative. One can only use coordinates adapted to several isometries for those isometries that generate an Abelian subgroup. However, even if the isometries do not commute, there are many conserved currents as isometries and this guarantees that there are as many combinations of the equations of motion as isometries that can be written as total derivatives. These combinations can be used to define on-shell duals of scalars. Carrying it out this program in the full theory can be, in practice, quite complicated.
@@ 5 A democratic pseudoaction for $\mathcal{N}=2B,d=10$ supergravity
The self-duality of the 5-form field strength forbids the existence of a covariant action and one must necessarily work with equations of motion. The equations of motion of this theory were first found in Ref. [33] in the Einstein frame and using a SU(1,1)/(1,1)/( 1 , 1 ) /U$(1)$ formulation of the coset space parametrized by the scalar fields. A pseudoaction which had to be supplemented by the self-duality constraint was first constructed in Ref. [27]. A pseudoaction containing the duals of the 1- and 3-form RR field strengths and some higher RR field strengths was constructed in Ref. [24]. While this action was “RR-democratic” it was certainly not “NSNS-democratic”. Furthermore, this incomplete democratization breaks the manifest SL$(2,\mathbb{R})$ symmetry of the theory. Our goal in this section is to improve on those results constructing a manifestly SL$(2,\mathbb{R})$-invariant and fully democratic pseudoaction using the results of the previous sections.
@@
$\mathbf{E}^{i}$ $=-d\left(\mathcal{M}^{ij}\star\mathcal{H}_{j}\right)-\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{j}\wedge\left(\mathcal{F}+\star\mathcal{F}\right)+\tfrac{1}{2}\varepsilon^{ij}\mathcal{B}_{j}\wedge\mathbf{E}\,,$ (5.15c)
@@ 5.1 Dualization of the 2-forms
$\mathbf{E}^{i}$ $=-\tfrac{1}{2}\left\{d\left(\mathcal{M}^{ij}\star\mathcal{H}_{j}\right)+\varepsilon^{ij}\mathcal{M}_{jk}\star\mathcal{H}^{k}\wedge\mathcal{F}\right\}+\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{j}\wedge(\star\mathcal{F}-\mathcal{F})$
$\hskip 14.22636pt+\tfrac{1}{6}\left(\varepsilon^{ij}\varepsilon^{mn}-4\varepsilon^{in}\varepsilon^{jm}+6\varepsilon^{im}\varepsilon^{jn}\right)\mathcal{B}_{m}\wedge\mathcal{H}_{n}\wedge(\mathcal{M}_{jk}\star\mathcal{H}^{k})-\tfrac{1}{4}\varepsilon^{im}\varepsilon^{jn}\mathcal{B}_{m}\wedge\mathcal{H}_{n}\wedge\mathcal{H}_{j}$
$\hskip 14.22636pt+\tfrac{1}{2}\varepsilon^{ij}\mathcal{B}_{j}\wedge\mathbf{E}+\tfrac{1}{6}\varepsilon^{ij}\varepsilon^{kl}\mathcal{B}_{i}\wedge\mathcal{B}_{j}\wedge\mathbf{E}_{k}\,,$ (5.24c)
$\mathbf{E}$ $=-\tfrac{1}{2}\left\{d\star\mathcal{F}-\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{i}\wedge\left(\mathcal{H}_{j}-2\mathcal{M}_{jk}\star\mathcal{H}^{k}\right)\right\}\,,$ (5.24d)
@@ 5.2 Dualization of the scalars
$\mathbf{E}^{i}$ $=-\tfrac{1}{2}\left\{d\left(\mathcal{M}^{ij}\star\mathcal{H}_{j}\right)+\varepsilon^{ij}\mathcal{M}_{jk}\star\mathcal{H}^{k}\wedge\mathcal{F}\right\}+\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{j}\wedge(\star\mathcal{F}-\mathcal{F})$
$\hskip 14.22636pt+\tfrac{1}{6}\left(\varepsilon^{ij}\varepsilon^{mn}-4\varepsilon^{in}\varepsilon^{jm}+6\varepsilon^{im}\varepsilon^{jn}\right)\mathcal{B}_{m}\wedge\mathcal{H}_{n}\wedge(\mathcal{M}_{jk}\star\mathcal{H}^{k})-\tfrac{1}{4}\varepsilon^{im}\varepsilon^{jn}\mathcal{B}_{m}\wedge\mathcal{H}_{n}\wedge\mathcal{H}_{j}$
$\hskip 14.22636pt+\tfrac{1}{2}\frac{\delta G_{A}}{\delta\mathcal{B}_{i}}\wedge\left[\mathfrak{M}^{AB}\star G_{B}-\hat{k}_{B}\right]+\tfrac{1}{2}\varepsilon^{ij}\mathcal{B}_{j}\wedge\mathbf{E}+\tfrac{1}{6}\varepsilon^{ij}\varepsilon^{kl}\mathcal{B}_{i}\wedge\mathcal{B}_{j}\wedge\mathbf{E}_{k}\,,$ (5.37c)
$\mathbf{E}$ $=-\tfrac{1}{2}\left\{d\star\mathcal{F}-\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{i}\wedge\left(\mathcal{H}_{j}-2\mathcal{M}_{jk}\star\mathcal{H}^{k}\right)\right\}+\tfrac{1}{2}\frac{\delta G_{A}}{\delta\mathcal{D}}\wedge\left[\mathfrak{M}^{AB}\star G_{B}-\hat{k}_{B}\right]\,,$ (5.37d)
@@ 6 Conclusions
    The 10-forms which are known to exist an play an important role in $\mathcal{N}=2B,d=10$ supergravity [39] should be added somehow to the pseudoaction in order to have a complete picture of the dualities between fields and fluxes.

### v3

@@ Introduction
The fact that an action for the higher-rank forms must also contain the lower-rank ones can be turned into an advantage if one manages to give consistence to the simultaneous presence of dual fields in the action. [^3] See Ref. [18] for the case of nonlinear electrodynamics and Ref. [19] for the extension to $(p+1)$-form potentials and their duals in arbitrary dimensions. The results of Ref. [19] could be used to formulate proper actions for the systems described in the current work via pseudoactions. For the free fields, this approach was introduced in Ref. [20] and discussed in detail for arbitrary $(p+1)$-form potentials and their duals in arbitrary dimensions in Ref. [21]. A solution to this problem is to use extensions of the Pasti-Sorokin-Tonin formalism [22] which is based on the introduction of an auxiliary scalar field in the action. This method has been used in Ref. [23, 24] [^4] See Ref. [25] and, specially, the more recent Ref. [26] for a review to construct covariant actions of $\mathcal{N}=2B,d=10$ supergravity containing the 4-form but also 8-form duals of the scalar fields and also in Refs. [27, 28] to construct an action of $\mathcal{D}=1,d=11$ supergravity containing the 3-form and its dual 6-form simultaneously. By dimensional reduction one can obtain an action of $\mathcal{N}=2B,d=10$ supergravity containing the fundamental and dual fields [29]. In a slightly different context, it has been used in Ref. [30] to construct a covariant worldvolume action of the M5-brane.
An alternative solution, proposed in Ref. [31], consists in including all the fields and treating them all on an equal footing as independent (making it democratic). This procedure doubles the degrees of freedom and one has to impose by hand the twisted duality [32, 33] relations between $(p+1)$- and $(d-p-3)$-forms only after the equations of motion have been derived from the action. Since the duality relations are not derived from the action, one is actually dealing with a pseudoaction. The pseudoaction introduced in Ref. [34] for $\mathcal{N}=2B,d=10$ supergravity, which includes the 4-form with selfdual 5-form field strength provides a good example: it contains an unconstrained (not selfdual) 4-form which describes twice the degrees of freedom of the selfdual one and the selfduality constraint must be imposed after the equations of motion have been derived.
@@
When the theory contains several scalar fields which parametrize a non-linear $\sigma$-model, things become more complicated. The shift symmetries are isometries of the $\sigma$-model metric. One can always use coordinates (scalar fields) adapted to a given isometry. In those coordinates the scalar field shifted by the isometry does not occur explicitly in the metric and the action can always be written in terms of its derivatives only. The equation of motion will be a total derivative. One can only use coordinates adapted to several isometries for those isometries that generate an Abelian subgroup. However, even if the isometries do not commute, there are many conserved currents as isometries and this guarantees that there are as many combinations of the equations of motion as isometries that can be written as total derivatives. These combinations can be used to define on-shell duals of scalars. Carrying it out this program in the full theory can be, in practice, quite complicated. See, for example, Refs. [36, 37, 38].
@@ 5 A democratic pseudoaction for $\mathcal{N}=2B,d=10$ supergravity
The self-duality of the 5-form field strength forbids the existence of a covariant action free of auxiliary fields and one, if one does not want to deal with auxiliary fields, one must necessarily work with equations of motion. The equations of motion of this theory were first found in Ref. [43] in the Einstein frame and using a SU(1,1)/(1,1)/( 1 , 1 ) /U$(1)$ formulation of the coset space parametrized by the scalar fields. A pseudoaction which had to be supplemented by the self-duality constraint was first constructed in Ref. [34]. A pseudoaction containing the duals of the 1- and 3-form RR field strengths and some higher RR field strengths was constructed in Ref. [31]. While this action was “RR-democratic” it was certainly not “NSNS-democratic”. Furthermore, this incomplete democratization breaks the manifest SL$(2,\mathbb{R})$ symmetry of the theory. Our goal in this section is to improve on those results constructing a manifestly SL$(2,\mathbb{R})$-invariant and fully democratic pseudoaction using the results of the previous sections.
@@
$\mathbf{E}^{i}$ =$=-\left\{d\left(\mathcal{M}^{ij}\star\mathcal{H}_{j}\right)+\varepsilon^{ij}\mathcal{H}_{j}\wedge\mathcal{F}\right\}+\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{j}\wedge\left(\mathcal{F}-\star\mathcal{F}\right)+\tfrac{1}{2}\varepsilon^{ij}\mathcal{B}_{j}\wedge\mathbf{E}\,,$ (5.15c)
@@ 5.1 Dualization of the 2-forms
$\mathbf{E}^{i}$ $=-\tfrac{1}{2}\left\{d\left(\mathcal{M}^{ij}\star\mathcal{H}_{j}\right)+\varepsilon^{ij}\mathcal{H}_{j}\wedge\mathcal{F}\right\}+\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{j}\wedge(\star\mathcal{F}-\mathcal{F})$
$\hskip 14.22636pt+\tfrac{1}{2}\varepsilon^{ij}\left(\mathcal{H}_{j}-\mathcal{M}_{jk}\star\mathcal{H}^{k}\right)\wedge\mathcal{F}+\tfrac{1}{2}\varepsilon^{ij}\mathcal{B}_{j}\wedge\mathbf{E}-\tfrac{2}{3}\varepsilon^{ij}\varepsilon^{kl}\mathcal{B}_{j}\wedge\mathcal{B}_{k}\wedge\mathbf{E}_{l}\,,$ (5.24c)
$\mathbf{E}$ $=-\tfrac{1}{2}\left\{d\star\mathcal{F}-\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{i}\wedge\left(\mathcal{H}_{j}-2\mathcal{M}_{jk}\star\mathcal{H}^{k}\right)\right\}+\varepsilon^{ij}\mathcal{B}_{i}\wedge\mathbf{E}_{j}\,,$ (5.24d)
@@ 5.2 Dualization of the scalars
$\mathbf{E}^{i}$ $=-\tfrac{1}{2}\left\{d\left(\mathcal{M}^{ij}\star\mathcal{H}_{j}\right)+\varepsilon^{ij}\mathcal{H}_{j}\wedge\mathcal{F}\right\}+\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{j}\wedge(\star\mathcal{F}-\mathcal{F})$
$\hskip 14.22636pt+\tfrac{1}{2}\varepsilon^{ij}\left(\mathcal{H}_{j}-\mathcal{M}_{jk}\star\mathcal{H}^{k}\right)\wedge\mathcal{F}+\tfrac{1}{2}\frac{\delta G_{A}}{\delta\mathcal{B}_{i}}\wedge\left[\mathfrak{M}^{AB}\star G_{B}-\hat{k}_{B}\right]$
$\hskip 14.22636pt+\tfrac{1}{2}\varepsilon^{ij}\mathcal{B}_{j}\wedge\mathbf{E}-\tfrac{2}{3}\varepsilon^{ij}\varepsilon^{kl}\mathcal{B}_{j}\wedge\mathcal{B}_{k}\wedge\mathbf{E}_{l}\,,$ (5.37c)
$\mathbf{E}$ $=-\tfrac{1}{2}\left\{d\star\mathcal{F}-\tfrac{1}{2}\varepsilon^{ij}\mathcal{H}_{i}\wedge\left(\mathcal{H}_{j}-2\mathcal{M}_{jk}\star\mathcal{H}^{k}\right)\right\}+\tfrac{1}{2}\frac{\delta G_{A}}{\delta\mathcal{D}}\wedge\left[\mathfrak{M}^{AB}\star G_{B}-\hat{k}_{B}\right]$
$\hskip 14.22636pt+\varepsilon^{ij}\mathcal{B}_{i}\wedge\mathbf{E}_{j}\,,$ (5.37d)
@@ 6 Conclusions
    The 10-forms which are known to exist and play an important role in $\mathcal{N}=2B,d=10$ supergravity [49] should be added somehow to the pseudoaction in order to have a complete picture of the dualities between fields and fluxes.